Mrs. Franklin, who is in the 37 percent tax bracket, owns a residential apartment building that generates $100,000 annual taxTax savings from this income-shifting arrangement
Answer #1

Answer :

As Franklin transferred 15% equity interest in building to two children
Reduction in her income = 100000*15%*2
= 30000
a Tax savings = 37%*30000
= 11100
b Tax liability for two Children's = 30000*12%
= 3600
c Net savings (a -b) = $7500